Describe the preparation of 4.70 kg of 12.0% (w/w) aqueous glycerol.
Answer #1

Total mass of solution ( water + glycerol) = 4.70 kg.

Since %(w/w) = (mass of solute/ mass of solution)x100%

Therefore, mass of glycerol required = (12/100)x4.70 kg = 0.564 kg.

Therefore, mass of glycerol 0.564 kg to be mixed with (4.7-0.564 = 4.136 kg) water to form a 12.0%(w/w) aqueous glycerol solution.